Is My Site Affected?

204 Resolved Issues:

Filter issues: resolved | all | current

Resolved – Global Issue: Cloud Backups
Resolved at 13:06 UTC on 17 Nov 2017 (1 day ago)
First posted at 11:59 UTC on 16 Nov 2017 and last updated at 13:06 UTC on 17 Nov 2017.
Resolved – Global Issue: Issue adding cards in the Bytemark Panel
Resolved at 13:08 UTC on 16 Nov 2017 (2 days ago)
First posted at 11:15 UTC on 16 Nov 2017 and last updated at 13:08 UTC on 16 Nov 2017.
Resolved – Global Issue: Intermittent network for VMs on Head14
Resolved at 13:33 UTC on 13 Nov 2017 (5 days ago)
First posted at 13:04 UTC on 13 Nov 2017 and last updated at 13:33 UTC on 13 Nov 2017.
Resolved – BigV - bigv1-head2: Some Cloud Servers in Manchester may be down.
Resolved at 13:19 UTC on 12 Nov 2017 (6 days ago)
First posted at 12:55 UTC on 12 Nov 2017 and last updated at 13:19 UTC on 12 Nov 2017.
This was a scheduled issue; it began at 12:45 UTC on 12 Nov 2017.
Resolved – Global Issue: head14 outage from DDoS
Resolved at 09:33 UTC on 10 Nov 2017 (8 days ago)
First posted at 09:30 UTC on 10 Nov 2017 and last updated at 09:33 UTC on 10 Nov 2017.
Resolved – Global Issue: Bytemark Panel
Resolved at 14:52 UTC on 01 Nov 2017 (17 days ago)
First posted at 13:32 UTC on 01 Nov 2017 and last updated at 14:52 UTC on 01 Nov 2017.
Resolved – Global Issue: Panel unable to start or stop VMs
Resolved at 23:15 UTC on 28 Oct 2017 (21 days ago)
First posted at 22:55 UTC on 28 Oct 2017 and last updated at 23:15 UTC on 28 Oct 2017.
Resolved – Global Issue: Bytemark Cloud disks
Resolved at 10:03 UTC on 26 Oct 2017 (23 days ago)
First posted at 09:59 UTC on 26 Oct 2017 and last updated at 10:03 UTC on 26 Oct 2017.
Resolved – Global Issue: Backup server Foxtrot maintenance
Resolved at 10:53 UTC on 25 Oct 2017 (24 days ago)
First posted at 13:45 UTC on 17 Oct 2017 and last updated at 10:53 UTC on 25 Oct 2017.
This was a scheduled issue; it began at 09:00 UTC on 24 Oct 2017.
Resolved – Global Issue: DDoS attack affecting Bytemark Cloud in Manchester
Resolved at 21:46 UTC on 22 Oct 2017 (27 days ago)
First posted at 18:27 UTC on 22 Oct 2017 and last updated at 21:46 UTC on 22 Oct 2017.
Resolved – Global Issue: LONAP peering withdrawn temporarily
Resolved at 19:40 UTC on 19 Oct 2017 (30 days ago)
First posted at 11:17 UTC on 18 Oct 2017 and last updated at 19:40 UTC on 19 Oct 2017.
Resolved – Global Issue: Temporary issue with Bytemark Panel - Cloud Server access and changes not operational
Resolved at 11:26 UTC on 04 Oct 2017 (about 2 months ago)
First posted at 11:23 UTC on 04 Oct 2017 and last updated at 11:26 UTC on 04 Oct 2017.
This was a scheduled issue; it began at 11:00 UTC on 04 Oct 2017.
Resolved – Global Issue: VNC connections in the panel are not available
Resolved at 09:52 UTC on 01 Oct 2017 (about 2 months ago)
First posted at 09:31 UTC on 01 Oct 2017 and last updated at 09:52 UTC on 01 Oct 2017.
Resolved – BigV - bigv1-tail21: Tail 21 unexpected reboot
Resolved at 06:42 UTC on 25 Sep 2017 (about 2 months ago)
First posted at 06:39 UTC on 25 Sep 2017 and last updated at 06:42 UTC on 25 Sep 2017.
Resolved – Global Issue: At Risk: Core Router Relocation, Manchester
Resolved at 19:58 UTC on 22 Sep 2017 (about 2 months ago)
First posted at 21:33 UTC on 20 Sep 2017 and last updated at 19:58 UTC on 22 Sep 2017.
This was a scheduled issue; it began at 06:30 UTC on 21 Sep 2017.
Resolved – Global Issue: At Risk: Core router fan tray replacement, Manchester
Resolved at 19:04 UTC on 17 Aug 2017 (3 months ago)
First posted at 12:01 UTC on 17 Aug 2017 and last updated at 19:04 UTC on 17 Aug 2017.
Resolved – Global Issue: At Risk: fibre maintenance in Manchester
Resolved at 10:26 UTC on 15 Aug 2017 (3 months ago)
First posted at 20:42 UTC on 14 Aug 2017 and last updated at 10:26 UTC on 15 Aug 2017.
This was a scheduled issue; it began at 17:30 UTC on 14 Aug 2017.
Resolved – Meta Issue: Short outage of Bytemark Panel and Bytemark Phone Lines
Resolved at 15:05 UTC on 14 Aug 2017 (3 months ago)
First posted at 14:40 UTC on 14 Aug 2017 and last updated at 15:05 UTC on 14 Aug 2017.
This was a scheduled issue; it began at 14:30 UTC on 14 Aug 2017.
Resolved – Global Issue: At Risk: remedial work in Leeds
Resolved at 22:05 UTC on 17 Jul 2017 (4 months ago)
First posted at 20:30 UTC on 17 Jul 2017 and last updated at 22:05 UTC on 17 Jul 2017.
This was a scheduled issue; it began at 21:00 UTC on 17 Jul 2017.
Resolved – Meta Issue: Symbiosis Documentation
Resolved at 18:45 UTC on 13 Jul 2017 (4 months ago)
First posted at 16:28 UTC on 13 Jul 2017 and last updated at 18:45 UTC on 13 Jul 2017.
Resolved – Global Issue: Bytemark Panel:
Resolved at 14:23 UTC on 11 Jul 2017 (4 months ago)
First posted at 14:18 UTC on 11 Jul 2017 and last updated at 14:23 UTC on 11 Jul 2017.
Resolved – Meta Issue: Forum downtime
Resolved at 14:22 UTC on 11 Jul 2017 (4 months ago)
First posted at 11:33 UTC on 11 Jul 2017 and last updated at 14:22 UTC on 11 Jul 2017.
This was a scheduled issue; it began at 14:00 UTC on 11 Jul 2017.
Resolved – Global Issue: At-Risk: reboot & re-location of router in York
Resolved at 23:57 UTC on 10 Jul 2017 (4 months ago)
First posted at 20:36 UTC on 10 Jul 2017 and last updated at 23:57 UTC on 10 Jul 2017.
This was a scheduled issue; it began at 21:00 UTC on 10 Jul 2017.
Resolved – Dedicated - fw1.yo26.yrk: Firewall problem
Resolved at 06:34 UTC on 07 Jul 2017 (4 months ago)
First posted at 06:01 UTC on 07 Jul 2017 and last updated at 06:34 UTC on 07 Jul 2017.
Resolved – Global Issue: DNS resolution for subdomains of bytemark.co.uk may be unavailable
Resolved at 11:10 UTC on 04 Jul 2017 (5 months ago)
First posted at 11:01 UTC on 04 Jul 2017 and last updated at 11:10 UTC on 04 Jul 2017.
This was a scheduled issue; it began at 10:30 UTC on 04 Jul 2017.
Resolved – Meta Issue: Management Panel unavailable
Resolved at 09:29 UTC on 04 Jul 2017 (5 months ago)
First posted at 07:53 UTC on 04 Jul 2017 and last updated at 09:29 UTC on 04 Jul 2017.
This was a scheduled issue; it began at 04:50 UTC on 04 Jul 2017.
Resolved – BigV - bigv1-tail56: Some Cloud Servers may have problems writing to disks.
Resolved at 21:56 UTC on 29 Jun 2017 (5 months ago)
First posted at 13:12 UTC on 29 Jun 2017 and last updated at 21:56 UTC on 29 Jun 2017.
Resolved – BigV - bigv1-brain1: Customers are not able to control Cloud Servers via panel.bytemark.co.uk
Resolved at 15:50 UTC on 29 Jun 2017 (5 months ago)
First posted at 15:29 UTC on 29 Jun 2017 and last updated at 15:50 UTC on 29 Jun 2017.
Resolved – VirtualMachineHost - host03.vmh: Legacy Virtual Servers on host03.vmh are currently down.
Resolved at 11:02 UTC on 06 Jun 2017 (6 months ago)
First posted at 08:37 UTC on 06 Jun 2017 and last updated at 11:02 UTC on 06 Jun 2017.
Resolved – Global Issue: Bytemark Cloud: Imaging Machines with an SSH Key
Resolved at 12:48 UTC on 02 Jun 2017 (6 months ago)
First posted at 15:15 UTC on 31 May 2017 and last updated at 12:48 UTC on 02 Jun 2017.
Resolved – Meta Issue: Some Bytemark sites are offline
Resolved at 15:39 UTC on 25 May 2017 (6 months ago)
First posted at 11:03 UTC on 25 May 2017 and last updated at 15:39 UTC on 25 May 2017.
Resolved – BigV - bigv1-head10: Cloud servers hosted on 'head10' down
Resolved at 13:58 UTC on 22 May 2017 (6 months ago)
First posted at 13:49 UTC on 22 May 2017 and last updated at 13:58 UTC on 22 May 2017.
Resolved – Meta Issue: around 120 vms died mid migration
Resolved at 12:56 UTC on 18 May 2017 (6 months ago)
First posted at 12:20 UTC on 18 May 2017 and last updated at 12:56 UTC on 18 May 2017.
Resolved – Meta Issue: Scheduled maintenance: Ticketing system (for responding to your emails)
Resolved at 18:12 UTC on 17 May 2017 (6 months ago)
First posted at 12:50 UTC on 12 May 2017 and last updated at 18:12 UTC on 17 May 2017.
This was a scheduled issue; it began at 16:00 UTC on 17 May 2017.
Resolved – Meta Issue: Issues with Bytemark telephone support
Resolved at 11:40 UTC on 17 May 2017 (6 months ago)
First posted at 10:49 UTC on 17 May 2017 and last updated at 11:40 UTC on 17 May 2017.
Resolved – BigV - bigv1-head54: Cloud Servers hosted on 'head54' down.
Resolved at 15:12 UTC on 15 May 2017 (6 months ago)
First posted at 14:50 UTC on 15 May 2017 and last updated at 15:12 UTC on 15 May 2017.
Resolved – Global Issue: At Risk: fibre cut affecting direct York to London paths
Resolved at 17:08 UTC on 14 May 2017 (6 months ago)
First posted at 14:29 UTC on 14 May 2017 and last updated at 17:08 UTC on 14 May 2017.
Resolved – Global Issue: BigV VM unreachability
Resolved at 08:56 UTC on 10 May 2017 (6 months ago)
First posted at 16:37 UTC on 08 May 2017 and last updated at 08:56 UTC on 10 May 2017.
Resolved – Global Issue: Bytemark Panel hanging
Resolved at 19:05 UTC on 08 May 2017 (6 months ago)
First posted at 14:51 UTC on 08 May 2017 and last updated at 19:05 UTC on 08 May 2017.
Resolved – Global Issue: At Risk: Software upgrades to cr2.lds, Leeds
Resolved at 07:20 UTC on 03 May 2017 (7 months ago)
First posted at 05:39 UTC on 03 May 2017 and last updated at 07:20 UTC on 03 May 2017.
Resolved – Global Issue: Ability to add new credit cards
Resolved at 09:04 UTC on 27 Apr 2017 (7 months ago)
First posted at 08:28 UTC on 27 Apr 2017 and last updated at 09:04 UTC on 27 Apr 2017.
Resolved – Global Issue: Servers down on tail21
Resolved at 01:37 UTC on 23 Apr 2017 (7 months ago)
First posted at 10:48 UTC on 22 Apr 2017 and last updated at 01:37 UTC on 23 Apr 2017.
Resolved – Global Issue: Read only pool on tail21
Resolved at 17:59 UTC on 19 Apr 2017 (7 months ago)
First posted at 16:52 UTC on 19 Apr 2017 and last updated at 17:59 UTC on 19 Apr 2017.
Resolved – Global Issue: At Risk: Software upgrades to cr4.man, Manchester
Resolved at 00:56 UTC on 12 Apr 2017 (7 months ago)
First posted at 22:07 UTC on 11 Apr 2017 and last updated at 00:56 UTC on 12 Apr 2017.
Resolved – Meta Issue: Bytemark web site is offline
Resolved at 14:04 UTC on 11 Apr 2017 (7 months ago)
First posted at 13:58 UTC on 11 Apr 2017 and last updated at 14:04 UTC on 11 Apr 2017.
Resolved – Global Issue: Email problems with Symbiosis this morning
Resolved at 12:38 UTC on 11 Apr 2017 (7 months ago)
First posted at 08:15 UTC on 11 Apr 2017 and last updated at 12:38 UTC on 11 Apr 2017.
Resolved – Global Issue: Brief interruption in network connectivity
Resolved at 11:54 UTC on 11 Apr 2017 (7 months ago)
First posted at 08:07 UTC on 11 Apr 2017 and last updated at 11:54 UTC on 11 Apr 2017.
Resolved – Global Issue: Cloud Servers being investigated and restarted
Resolved at 10:45 UTC on 11 Apr 2017 (7 months ago)
First posted at 09:33 UTC on 11 Apr 2017 and last updated at 10:45 UTC on 11 Apr 2017.
Resolved – Global Issue: Problems with our ticketing system
Resolved at 09:54 UTC on 10 Apr 2017 (7 months ago)
First posted at 09:50 UTC on 10 Apr 2017 and last updated at 09:54 UTC on 10 Apr 2017.
Resolved – Global Issue: Potentially Poor IO on BigV
Resolved at 09:48 UTC on 05 Apr 2017 (8 months ago)
First posted at 11:04 UTC on 03 Apr 2017 and last updated at 09:48 UTC on 05 Apr 2017.
Resolved – Global Issue: [AT RISK] York DC Network Connectivity
Resolved at 13:48 UTC on 28 Mar 2017 (8 months ago)
First posted at 09:29 UTC on 28 Mar 2017 and last updated at 13:48 UTC on 28 Mar 2017.
This was a scheduled issue; it began at 08:00 UTC on 28 Mar 2017.
Resolved – Meta Issue: Power outage in one of our (York) racks
Resolved at 11:22 UTC on 23 Mar 2017 (8 months ago)
First posted at 08:31 UTC on 23 Mar 2017 and last updated at 11:22 UTC on 23 Mar 2017.
Resolved – Dedicated - mirror: Scheduled maintenance mirror.bytemark.co.uk Thursday 23rd March 2017
Resolved at 10:52 UTC on 23 Mar 2017 (8 months ago)
First posted at 13:05 UTC on 21 Mar 2017 and last updated at 10:52 UTC on 23 Mar 2017.
This was a scheduled issue; it began at 08:00 UTC on 23 Mar 2017.
Resolved – Meta Issue: Potential failure in inbound or outbound calls
Resolved at 17:18 UTC on 20 Mar 2017 (8 months ago)
First posted at 16:49 UTC on 20 Mar 2017 and last updated at 17:18 UTC on 20 Mar 2017.
Resolved – Global Issue: upload.ns.bytemark.co.uk is down
Resolved at 07:53 UTC on 14 Mar 2017 (8 months ago)
First posted at 06:25 UTC on 13 Mar 2017 and last updated at 07:53 UTC on 14 Mar 2017.
Resolved – Switch - as1.5.dc1.yo26.yrk: PDU and switch failure in rack 05.dc1.yo26.yrk
Resolved at 13:25 UTC on 06 Mar 2017 (9 months ago)
First posted at 13:19 UTC on 06 Mar 2017 and last updated at 13:25 UTC on 06 Mar 2017.
Resolved – Meta Issue: LINX LON1 Maintenance
Resolved at 11:48 UTC on 02 Mar 2017 (9 months ago)
First posted at 23:49 UTC on 01 Mar 2017 and last updated at 11:48 UTC on 02 Mar 2017.
This was a scheduled issue; it began at 00:00 UTC on 02 Mar 2017.
Resolved – Global Issue: At Risk: York DC network
Resolved at 14:48 UTC on 01 Mar 2017 (9 months ago)
First posted at 14:57 UTC on 28 Feb 2017 and last updated at 14:48 UTC on 01 Mar 2017.
This was a scheduled issue; it began at 13:00 UTC on 01 Mar 2017.
Resolved – Global Issue: Central database cluster malfunction
Resolved at 15:19 UTC on 15 Feb 2017 (9 months ago)
First posted at 20:57 UTC on 09 Feb 2017 and last updated at 15:19 UTC on 15 Feb 2017.
This was a scheduled issue; it began at 20:00 UTC on 09 Feb 2017.
Resolved – Global Issue: Unexpected reboot of Gateway router in York
Resolved at 18:06 UTC on 06 Feb 2017 (9 months ago)
First posted at 17:02 UTC on 06 Feb 2017 and last updated at 18:06 UTC on 06 Feb 2017.
Resolved – Global Issue: York DC Core Network Loop
Resolved at 15:25 UTC on 02 Feb 2017 (10 months ago)
First posted at 16:23 UTC on 25 Jan 2017 and last updated at 15:25 UTC on 02 Feb 2017.
Resolved – Global Issue: Core router upgrade, Manchester
Resolved at 17:06 UTC on 25 Jan 2017 (10 months ago)
First posted at 18:27 UTC on 23 Jan 2017 and last updated at 17:06 UTC on 25 Jan 2017.
This was a scheduled issue; it began at 22:30 UTC on 24 Jan 2017.
Resolved – BigV - bigv1-head60: Head 60 required reboot
Resolved at 10:54 UTC on 11 Jan 2017 (10 months ago)
First posted at 23:04 UTC on 08 Jan 2017 and last updated at 10:54 UTC on 11 Jan 2017.
Resolved – Issue 148: Serial console access for some Cloud Servers is temporarily available
Resolved at 21:50 UTC on 16 Dec 2016 (11 months ago)
First posted at 21:26 UTC on 16 Dec 2016 and last updated at 21:50 UTC on 16 Dec 2016.
Resolved – Global Issue: Backup Server - Foxtrot RAID rebuilding
Resolved at 10:56 UTC on 16 Dec 2016 (11 months ago)
First posted at 11:35 UTC on 05 Dec 2016 and last updated at 10:56 UTC on 16 Dec 2016.
Resolved – BigV - bigv1-tail19: Potential slowdown on Tail 19
Resolved at 14:20 UTC on 29 Nov 2016 (12 months ago)
First posted at 12:46 UTC on 26 Nov 2016 and last updated at 14:20 UTC on 29 Nov 2016.
Resolved – Global Issue: Issues with our login/authentication service.
Resolved at 14:02 UTC on 24 Nov 2016 (12 months ago)
First posted at 12:54 UTC on 24 Nov 2016 and last updated at 14:02 UTC on 24 Nov 2016.
Resolved – BigV - bigv1-head13: Loss of networking to VMs on head13
Resolved at 11:14 UTC on 22 Nov 2016 (12 months ago)
First posted at 10:37 UTC on 18 Nov 2016 and last updated at 11:14 UTC on 22 Nov 2016.
Resolved – Meta Issue: Planned upgrade/migration of upload.ns.bytemark.co.uk.
Resolved at 08:28 UTC on 22 Nov 2016 (12 months ago)
First posted at 06:57 UTC on 19 Nov 2016 and last updated at 08:28 UTC on 22 Nov 2016.
This was a scheduled issue; it began at 05:00 UTC on 22 Nov 2016.
Resolved – Global Issue: Core router replacement, York
Resolved at 03:35 UTC on 04 Nov 2016 (about 1 year ago)
First posted at 21:39 UTC on 03 Nov 2016 and last updated at 03:35 UTC on 04 Nov 2016.
This was a scheduled issue; it began at 22:00 UTC on 03 Nov 2016.
Resolved – BigV - bigv1-head52: BigV uk0 outage
Resolved at 03:26 UTC on 04 Nov 2016 (about 1 year ago)
First posted at 02:15 UTC on 04 Nov 2016 and last updated at 03:26 UTC on 04 Nov 2016.
Resolved – BigV - bigv1-head23: Disk issue on cloud head
Resolved at 22:29 UTC on 29 Oct 2016 (about 1 year ago)
First posted at 21:41 UTC on 29 Oct 2016 and last updated at 22:29 UTC on 29 Oct 2016.
Resolved – BigV - bigv1-head10: BigV Head10 Issue
Resolved at 10:43 UTC on 27 Oct 2016 (about 1 year ago)
First posted at 08:35 UTC on 27 Oct 2016 and last updated at 10:43 UTC on 27 Oct 2016.
Resolved – Switch - as1.22.dc1.yo26.yrk: Unexpected reboot of as1.22
Resolved at 15:15 UTC on 25 Oct 2016 (about 1 year ago)
First posted at 15:15 UTC on 25 Oct 2016.
Resolved – Global Issue: Core router replacement, York
Resolved at 04:38 UTC on 21 Oct 2016 (about 1 year ago)
First posted at 20:17 UTC on 19 Oct 2016 and last updated at 04:38 UTC on 21 Oct 2016.
This was a scheduled issue; it began at 19:00 UTC on 20 Oct 2016.
Resolved – VirtualMachineHost - host26.vmh: host26.vmh rebooting
Resolved at 18:09 UTC on 14 Oct 2016 (about 1 year ago)
First posted at 16:26 UTC on 14 Oct 2016 and last updated at 18:09 UTC on 14 Oct 2016.
Resolved – Global Issue: Slow I/O performance on some Cloud Servers
Resolved at 16:08 UTC on 12 Oct 2016 (about 1 year ago)
First posted at 12:24 UTC on 12 Oct 2016 and last updated at 16:08 UTC on 12 Oct 2016.
Resolved – Global Issue: BigV: Bytemark DB server outage
Resolved at 15:47 UTC on 06 Oct 2016 (about 1 year ago)
First posted at 14:48 UTC on 06 Oct 2016 and last updated at 15:47 UTC on 06 Oct 2016.
Resolved – Global Issue: Serial console access for some cloud servers is temporarily unavailable
Resolved at 11:53 UTC on 05 Oct 2016 (about 1 year ago)
First posted at 09:34 UTC on 05 Oct 2016 and last updated at 11:53 UTC on 05 Oct 2016.
Resolved – Global Issue: Dedicated host console server (York) currently unavailable
Resolved at 18:23 UTC on 30 Sep 2016 (about 1 year ago)
First posted at 18:03 UTC on 30 Sep 2016 and last updated at 18:23 UTC on 30 Sep 2016.
Resolved – Global Issue: Unexpected core router failure, York
Resolved at 16:13 UTC on 27 Sep 2016 (about 1 year ago)
First posted at 11:41 UTC on 27 Sep 2016 and last updated at 16:13 UTC on 27 Sep 2016.
Resolved – Global Issue: Cloud servers control problem (York)
Resolved at 14:48 UTC on 27 Sep 2016 (about 1 year ago)
First posted at 12:48 UTC on 27 Sep 2016 and last updated at 14:48 UTC on 27 Sep 2016.
Resolved – Global Issue: Networking issue (spanned services)
Resolved at 08:30 UTC on 15 Sep 2016 (about 1 year ago)
First posted at 08:03 UTC on 15 Sep 2016 and last updated at 08:30 UTC on 15 Sep 2016.
Resolved – Global Issue: Core switch relocation, York
Resolved at 17:13 UTC on 13 Sep 2016 (about 1 year ago)
First posted at 15:29 UTC on 30 Aug 2016 and last updated at 17:13 UTC on 13 Sep 2016.
This was a scheduled issue; it began at 05:00 UTC on 13 Sep 2016.
Resolved – Meta Issue: ContentDNS upload failures
Resolved at 18:36 UTC on 05 Sep 2016 (about 1 year ago)
First posted at 17:11 UTC on 05 Sep 2016 and last updated at 18:36 UTC on 05 Sep 2016.
This was a scheduled issue; it began at 13:00 UTC on 05 Sep 2016.
Resolved – Global Issue: Minor: Router Reload, London
Resolved at 02:02 UTC on 03 Sep 2016 (about 1 year ago)
First posted at 00:23 UTC on 03 Sep 2016 and last updated at 02:02 UTC on 03 Sep 2016.
Resolved – BigV - bigv1-head25: Cloud Server outage for customers on head25
Resolved at 15:14 UTC on 26 Aug 2016 (about 1 year ago)
First posted at 15:14 UTC on 26 Aug 2016.
Resolved – BigV - bigv1-head25: Cloud Server outage for customers on head25
Resolved at 21:30 UTC on 24 Aug 2016 (about 1 year ago)
First posted at 21:29 UTC on 24 Aug 2016 and last updated at 21:30 UTC on 24 Aug 2016.
Resolved – BigV - bigv1-brain1: Control issue affecting York heads
Resolved at 18:22 UTC on 23 Aug 2016 (about 1 year ago)
First posted at 16:15 UTC on 23 Aug 2016 and last updated at 18:22 UTC on 23 Aug 2016.
This was a scheduled issue; it began at 11:30 UTC on 23 Aug 2016.
Resolved – Global Issue: DoS attack against Manchester IP
Resolved at 13:57 UTC on 23 Aug 2016 (about 1 year ago)
First posted at 13:43 UTC on 23 Aug 2016 and last updated at 13:57 UTC on 23 Aug 2016.
Resolved – BigV - bigv1-head1: Repeat of disk issue on cloud head
Resolved at 18:31 UTC on 22 Aug 2016 (about 1 year ago)
First posted at 18:26 UTC on 22 Aug 2016 and last updated at 18:31 UTC on 22 Aug 2016.
Resolved – BigV - bigv1-head1: Disk issue on cloud head
Resolved at 03:49 UTC on 17 Aug 2016 (over 1 year ago)
First posted at 03:14 UTC on 17 Aug 2016 and last updated at 03:49 UTC on 17 Aug 2016.
Resolved – BigV - bigv1-head55: Disk I/O problem for cloud servers hosted on the affected head
Resolved at 22:57 UTC on 14 Aug 2016 (over 1 year ago)
First posted at 22:19 UTC on 14 Aug 2016 and last updated at 22:57 UTC on 14 Aug 2016.
Resolved – Global Issue: Maintenance to York DC Network
Resolved at 02:07 UTC on 11 Aug 2016 (over 1 year ago)
First posted at 16:57 UTC on 18 Jul 2016 and last updated at 02:07 UTC on 11 Aug 2016.
This was a scheduled issue; it began at 01:00 UTC on 18 Jul 2016.
Resolved – Global Issue: Bytemark Cloud head02 Spontaneous Reboot
Resolved at 10:51 UTC on 10 Aug 2016 (over 1 year ago)
First posted at 22:16 UTC on 05 Aug 2016 and last updated at 10:51 UTC on 10 Aug 2016.
This was a scheduled issue; it began at 21:08 UTC on 05 Aug 2016.
Resolved – Global Issue: Bytemark Cloud head01 Spontaneous Reboot
Resolved at 10:50 UTC on 10 Aug 2016 (over 1 year ago)
First posted at 22:04 UTC on 03 Aug 2016 and last updated at 10:50 UTC on 10 Aug 2016.
Resolved – Global Issue: Cloud Services Brain Currently Experiencing Problems
Resolved at 15:37 UTC on 05 Aug 2016 (over 1 year ago)
First posted at 01:13 UTC on 02 Aug 2016 and last updated at 15:37 UTC on 05 Aug 2016.
Resolved – BigV - bigv1-tail52: Slow IO on tail52
Resolved at 16:00 UTC on 26 Jul 2016 (over 1 year ago)
First posted at 10:07 UTC on 26 Jul 2016 and last updated at 16:00 UTC on 26 Jul 2016.
Resolved – Global Issue: Disruption to some Bytemark Cloud disks in York
Resolved at 11:22 UTC on 25 Jul 2016 (over 1 year ago)
First posted at 11:17 UTC on 25 Jul 2016 and last updated at 11:22 UTC on 25 Jul 2016.
Resolved – BigV - bigv1-tail5: Archive storage on tail5 has become partially read-only
Resolved at 23:54 UTC on 21 Jul 2016 (over 1 year ago)
First posted at 22:31 UTC on 21 Jul 2016 and last updated at 23:54 UTC on 21 Jul 2016.
Resolved – BigV - bigv1-head22: bigv1-head22 reboot
Resolved at 03:30 UTC on 24 Jun 2016 (over 1 year ago)
First posted at 02:33 UTC on 24 Jun 2016 and last updated at 03:30 UTC on 24 Jun 2016.
Resolved – BigV - bigv1-head24: BigV1-head24 reboot
Resolved at 02:52 UTC on 23 Jun 2016 (over 1 year ago)
First posted at 02:23 UTC on 23 Jun 2016 and last updated at 02:52 UTC on 23 Jun 2016.
Resolved – VirtualMachineHost - host11.vmh: Legacy Virtual Machine Host 11 Rebooted
Resolved at 19:25 UTC on 12 Jun 2016 (over 1 year ago)
First posted at 18:24 UTC on 12 Jun 2016 and last updated at 19:25 UTC on 12 Jun 2016.
Resolved – Meta Issue: Ongoing issues with backup server Waltz
Resolved at 10:47 UTC on 09 Jun 2016 (over 1 year ago)
First posted at 16:17 UTC on 29 Apr 2016 and last updated at 10:47 UTC on 09 Jun 2016.
Resolved – Global Issue: mirror.bytemark.co.uk
Resolved at 15:57 UTC on 17 May 2016 (over 1 year ago)
First posted at 15:41 UTC on 17 May 2016 and last updated at 15:57 UTC on 17 May 2016.
Resolved – Global Issue: Vendor Dark Fibre Maintenance, Manchester (At Risk)
Resolved at 09:18 UTC on 17 May 2016 (over 1 year ago)
First posted at 17:00 UTC on 16 May 2016 and last updated at 09:18 UTC on 17 May 2016.
This was a scheduled issue; it began at 13:30 UTC on 16 May 2016.
Resolved – Global Issue: BigV brain outage
Resolved at 14:02 UTC on 16 May 2016 (over 1 year ago)
First posted at 12:11 UTC on 16 May 2016 and last updated at 14:02 UTC on 16 May 2016.
Resolved – BigV - bigv1-tail51: Some BigV disks can't be reached at the moment
Resolved at 12:42 UTC on 13 May 2016 (over 1 year ago)
First posted at 11:46 UTC on 13 May 2016 and last updated at 12:42 UTC on 13 May 2016.
Resolved – BigV - bigv1-head24: BigV1 Head24 Reboot (again)
Resolved at 08:11 UTC on 13 May 2016 (over 1 year ago)
First posted at 02:28 UTC on 13 May 2016 and last updated at 08:11 UTC on 13 May 2016.
Resolved – BigV - bigv1-head24: Head 24 Restarted
Resolved at 10:27 UTC on 12 May 2016 (over 1 year ago)
First posted at 02:20 UTC on 12 May 2016 and last updated at 10:27 UTC on 12 May 2016.
Resolved – BigV - bigv1-head56: Disc I/O issue over night with BigV VM's on head56
Resolved at 09:59 UTC on 29 Apr 2016 (over 1 year ago)
First posted at 09:59 UTC on 29 Apr 2016.
Resolved – Global Issue: Firmware Upgrades on console servers
Resolved at 12:55 UTC on 27 Apr 2016 (over 1 year ago)
First posted at 09:47 UTC on 27 Apr 2016 and last updated at 12:55 UTC on 27 Apr 2016.
Resolved – Global Issue: Moving DNS Server
Resolved at 15:15 UTC on 26 Apr 2016 (over 1 year ago)
First posted at 13:04 UTC on 26 Apr 2016 and last updated at 15:15 UTC on 26 Apr 2016.
This was a scheduled issue; it began at 14:30 UTC on 26 Apr 2016.
Resolved – VirtualMachineHost - host22.vmh: Brief network outage for legacy VMs on host22
Resolved at 14:06 UTC on 26 Apr 2016 (over 1 year ago)
First posted at 11:45 UTC on 26 Apr 2016 and last updated at 14:06 UTC on 26 Apr 2016.
This was a scheduled issue; it began at 14:00 UTC on 26 Apr 2016.
Resolved – Meta Issue: Backup server Waltz RAID rebuilding, offline overnight.
Resolved at 08:57 UTC on 26 Apr 2016 (over 1 year ago)
First posted at 16:00 UTC on 25 Apr 2016 and last updated at 08:57 UTC on 26 Apr 2016.
This was a scheduled issue; it began at 16:05 UTC on 25 Apr 2016.
Resolved – Global Issue: Loss of connectivity to rack 6
Resolved at 12:54 UTC on 20 Apr 2016 (over 1 year ago)
First posted at 11:42 UTC on 20 Apr 2016 and last updated at 12:54 UTC on 20 Apr 2016.
Resolved – BigV - bigv1-brain1: BigV Brain Currently Not Responding
Resolved at 12:56 UTC on 15 Apr 2016 (over 1 year ago)
First posted at 10:40 UTC on 15 Apr 2016 and last updated at 12:56 UTC on 15 Apr 2016.
Resolved – Global Issue: Raw Power Feed Maintenace
Resolved at 16:36 UTC on 01 Apr 2016 (over 1 year ago)
First posted at 16:29 UTC on 01 Apr 2016 and last updated at 16:36 UTC on 01 Apr 2016.
This was a scheduled issue; it began at 07:30 UTC on 05 Apr 2016.
Resolved – PowerDistributionUnit - msA.22.dc1.yo26: Rack 22 power distribution unit A reboot
Resolved at 10:24 UTC on 01 Apr 2016 (over 1 year ago)
First posted at 09:15 UTC on 31 Mar 2016 and last updated at 10:24 UTC on 01 Apr 2016.
This was a scheduled issue; it began at 07:30 UTC on 01 Apr 2016.
Resolved – Meta Issue: BigV Brain Currently Not Responding
Resolved at 06:28 UTC on 01 Apr 2016 (over 1 year ago)
First posted at 05:34 UTC on 01 Apr 2016 and last updated at 06:28 UTC on 01 Apr 2016.
Resolved – Global Issue: Short network disruptions tonight (specific services)
Resolved at 22:28 UTC on 29 Mar 2016 (over 1 year ago)
First posted at 16:30 UTC on 29 Mar 2016 and last updated at 22:28 UTC on 29 Mar 2016.
This was a scheduled issue; it began at 17:00 UTC on 29 Mar 2016.
Resolved – PowerDistributionUnit - msA.15.dc1.yo26.yrk: Loss of power to rack 15
Resolved at 12:26 UTC on 28 Mar 2016 (over 1 year ago)
First posted at 10:27 UTC on 28 Mar 2016 and last updated at 12:26 UTC on 28 Mar 2016.
Resolved – Global Issue: Core switch relocation, York
Resolved at 19:22 UTC on 17 Mar 2016 (over 1 year ago)
First posted at 10:52 UTC on 14 Mar 2016 and last updated at 19:22 UTC on 17 Mar 2016.
This was a scheduled issue; it began at 10:00 UTC on 17 Mar 2016.
Resolved – Meta Issue: Forum reconfiguration
Resolved at 13:16 UTC on 17 Mar 2016 (over 1 year ago)
First posted at 15:43 UTC on 16 Mar 2016 and last updated at 13:16 UTC on 17 Mar 2016.
This was a scheduled issue; it began at 13:00 UTC on 17 Mar 2016.
Resolved – Global Issue: Scheduled database maintenance for BigV
Resolved at 17:44 UTC on 16 Mar 2016 (over 1 year ago)
First posted at 14:51 UTC on 09 Mar 2016 and last updated at 17:44 UTC on 16 Mar 2016.
This was a scheduled issue; it began at 17:00 UTC on 16 Mar 2016.
Resolved – Global Issue: Waltz backup server unavailable
Resolved at 00:32 UTC on 15 Mar 2016 (over 1 year ago)
First posted at 09:53 UTC on 14 Mar 2016 and last updated at 00:32 UTC on 15 Mar 2016.
Resolved – Global Issue: IX Leeds Peering Interrupted Briefly
Resolved at 14:19 UTC on 10 Mar 2016 (over 1 year ago)
First posted at 14:19 UTC on 10 Mar 2016.
Resolved – Meta Issue: Forum scheduled down time.
Resolved at 13:42 UTC on 01 Mar 2016 (over 1 year ago)
First posted at 15:50 UTC on 29 Feb 2016 and last updated at 13:42 UTC on 01 Mar 2016.
This was a scheduled issue; it began at 13:00 UTC on 01 Mar 2016.
Resolved – Global Issue: Server moves within our York Data Centre
Resolved at 07:18 UTC on 24 Feb 2016 (over 1 year ago)
First posted at 13:55 UTC on 23 Feb 2016 and last updated at 07:18 UTC on 24 Feb 2016.
This was a scheduled issue; it began at 05:00 UTC on 24 Feb 2016.
Resolved – Global Issue: [Planned] Switch upgrades in Manchester
Resolved at 05:38 UTC on 24 Feb 2016 (over 1 year ago)
First posted at 10:37 UTC on 23 Feb 2016 and last updated at 05:38 UTC on 24 Feb 2016.
This was a scheduled issue; it began at 05:00 UTC on 24 Feb 2016.
Resolved – Meta Issue: Delayed reboots this morning
Resolved at 12:45 UTC on 21 Feb 2016 (over 1 year ago)
First posted at 12:45 UTC on 21 Feb 2016.
Resolved – Meta Issue: LINX Port Upgrade
Resolved at 15:17 UTC on 10 Feb 2016 (almost 2 years ago)
First posted at 14:12 UTC on 08 Feb 2016 and last updated at 15:17 UTC on 10 Feb 2016.
This was a scheduled issue; it began at 15:00 UTC on 08 Feb 2016.
Resolved – BigV - bigvint-head12: Various Bytemark services down due to head failure on internal BigV cluster
Resolved at 08:29 UTC on 10 Feb 2016 (almost 2 years ago)
First posted at 08:15 UTC on 10 Feb 2016 and last updated at 08:29 UTC on 10 Feb 2016.
Resolved – Global Issue: Phone System Maintenance
Resolved at 13:36 UTC on 08 Feb 2016 (almost 2 years ago)
First posted at 12:05 UTC on 08 Feb 2016 and last updated at 13:36 UTC on 08 Feb 2016.
This was a scheduled issue; it began at 13:30 UTC on 08 Feb 2016.
Resolved – Global Issue: Phones not working
Resolved at 11:04 UTC on 08 Feb 2016 (almost 2 years ago)
First posted at 09:49 UTC on 08 Feb 2016 and last updated at 11:04 UTC on 08 Feb 2016.
Resolved – Global Issue: Temporary API failure for BigV in York
Resolved at 15:24 UTC on 31 Jan 2016 (almost 2 years ago)
First posted at 15:24 UTC on 31 Jan 2016.
Resolved – Router - fw4.man: Denial of service attack against Manchester customer
Resolved at 03:33 UTC on 30 Jan 2016 (almost 2 years ago)
First posted at 00:38 UTC on 30 Jan 2016 and last updated at 03:33 UTC on 30 Jan 2016.
Resolved – Global Issue: BigV: Slow archive performance on tail4
Resolved at 15:05 UTC on 29 Jan 2016 (almost 2 years ago)
First posted at 09:57 UTC on 29 Jan 2016 and last updated at 15:05 UTC on 29 Jan 2016.
Resolved – Global Issue: BigV brain down (panel and CLI operations are unavailable)
Resolved at 09:04 UTC on 26 Jan 2016 (almost 2 years ago)
First posted at 08:57 UTC on 26 Jan 2016 and last updated at 09:04 UTC on 26 Jan 2016.
Resolved – Global Issue: [Planned] Generator Maintenance
Resolved at 17:27 UTC on 25 Jan 2016 (almost 2 years ago)
First posted at 07:43 UTC on 25 Jan 2016 and last updated at 17:27 UTC on 25 Jan 2016.
This was a scheduled issue; it began at 09:00 UTC on 25 Jan 2016.
Resolved – Global Issue: [At Risk] Software upgrade to cs1.reynolds.man
Resolved at 03:01 UTC on 20 Jan 2016 (almost 2 years ago)
First posted at 18:07 UTC on 11 Jan 2016 and last updated at 03:01 UTC on 20 Jan 2016.
This was a scheduled issue; it began at 22:00 UTC on 19 Jan 2016.
Resolved – Router - fw2.man: Packet loss through fw2.man
Resolved at 15:47 UTC on 15 Jan 2016 (almost 2 years ago)
First posted at 13:55 UTC on 15 Jan 2016 and last updated at 15:47 UTC on 15 Jan 2016.
Resolved – Global Issue: Unexpected reboot on BigV Head 20
Resolved at 10:47 UTC on 07 Jan 2016 (almost 2 years ago)
First posted at 10:47 UTC on 07 Jan 2016.
Resolved – Meta Issue: Introduction of secondary core router, Leeds
Resolved at 15:59 UTC on 05 Jan 2016 (almost 2 years ago)
First posted at 14:07 UTC on 05 Jan 2016 and last updated at 15:59 UTC on 05 Jan 2016.
Resolved – BigV - bigv1-head1: Unexpected outage of a BigV host.
Resolved at 10:01 UTC on 22 Dec 2015 (almost 2 years ago)
First posted at 08:51 UTC on 22 Dec 2015 and last updated at 10:01 UTC on 22 Dec 2015.
Resolved – Router - fw4.man: Denial of service attack caused unavailability
Resolved at 14:38 UTC on 21 Dec 2015 (almost 2 years ago)
First posted at 14:38 UTC on 21 Dec 2015.
This was a scheduled issue; it began at 08:10 UTC on 15 Dec 2015.
Resolved – PowerDistributionUnit - msB.5.dc1.yo26.yrk: Rack 5 partial loss of power
Resolved at 13:39 UTC on 16 Dec 2015 (almost 2 years ago)
First posted at 13:36 UTC on 16 Dec 2015 and last updated at 13:39 UTC on 16 Dec 2015.
This was a scheduled issue; it began at 13:06 UTC on 16 Dec 2015.
Resolved – BigV - bigv1-tail51: Slow performance on tail51
Resolved at 08:42 UTC on 04 Dec 2015 (almost 2 years ago)
First posted at 07:22 UTC on 04 Dec 2015 and last updated at 08:42 UTC on 04 Dec 2015.
Resolved – Global Issue: Level 3 fibre joint work, York
Resolved at 06:43 UTC on 03 Dec 2015 (almost 2 years ago)
First posted at 20:49 UTC on 01 Dec 2015 and last updated at 06:43 UTC on 03 Dec 2015.
This was a scheduled issue; it began at 01:00 UTC on 03 Dec 2015.
Resolved – BigV - bigv1-brain2: BigV brain down
Resolved at 20:01 UTC on 02 Dec 2015 (almost 2 years ago)
First posted at 17:04 UTC on 02 Dec 2015 and last updated at 20:01 UTC on 02 Dec 2015.
Resolved – Global Issue: Replacement of Core Router, Manchester
Resolved at 12:24 UTC on 25 Nov 2015 (almost 2 years ago)
First posted at 19:01 UTC on 24 Nov 2015 and last updated at 12:24 UTC on 25 Nov 2015.
This was a scheduled issue; it began at 07:00 UTC on 25 Nov 2015.
Resolved – Global Issue: 'b.ns.bytemark.co.uk' unreachability
Resolved at 06:20 UTC on 24 Nov 2015 (almost 2 years ago)
First posted at 05:52 UTC on 24 Nov 2015 and last updated at 06:20 UTC on 24 Nov 2015.
Resolved – VirtualMachineHost - host02.vmh: Legacy VM host02 failure
Resolved at 11:36 UTC on 17 Nov 2015 (about 2 years ago)
First posted at 10:31 UTC on 17 Nov 2015 and last updated at 11:36 UTC on 17 Nov 2015.
Resolved – Meta Issue: Ticket System: Database Maintenance on Wednesday Evening
Resolved at 10:53 UTC on 12 Nov 2015 (about 2 years ago)
First posted at 13:47 UTC on 10 Nov 2015 and last updated at 10:53 UTC on 12 Nov 2015.
This was a scheduled issue; it began at 19:00 UTC on 11 Nov 2015.
Resolved – Global Issue: [planned] [at risk] Repatching York Switch uplinks
Resolved at 13:43 UTC on 19 Oct 2015 (about 2 years ago)
First posted at 12:05 UTC on 19 Oct 2015 and last updated at 13:43 UTC on 19 Oct 2015.
This was a scheduled issue; it began at 12:30 UTC on 19 Oct 2015.
Resolved – Global Issue: IX Manchester peering, shifting to new router
Resolved at 15:14 UTC on 15 Oct 2015 (about 2 years ago)
First posted at 13:09 UTC on 15 Oct 2015 and last updated at 15:14 UTC on 15 Oct 2015.
Resolved – Meta Issue: Status Page will be down briefly on Thursday, 15 Oct at 2:00 pm BST for an update
Resolved at 13:26 UTC on 15 Oct 2015 (about 2 years ago)
First posted at 14:00 UTC on 14 Oct 2015 and last updated at 13:26 UTC on 15 Oct 2015.
Resolved – Global Issue: [Planned] [At risk] Repatching York Switch uplinks
Resolved at 15:58 UTC on 14 Oct 2015 (about 2 years ago)
First posted at 11:53 UTC on 13 Oct 2015 and last updated at 15:58 UTC on 14 Oct 2015.
Resolved – Global Issue: BigV head19 failure
Resolved at 06:59 UTC on 14 Oct 2015 (about 2 years ago)
First posted at 06:26 UTC on 14 Oct 2015 and last updated at 06:59 UTC on 14 Oct 2015.
Resolved – Global Issue: Bigv Head Failure
Resolved at 05:59 UTC on 02 Oct 2015 (about 2 years ago)
First posted at 04:53 UTC on 02 Oct 2015 and last updated at 05:59 UTC on 02 Oct 2015.
Resolved – Global Issue: Re-routing network link between Manchester & Leeds
Resolved at 17:40 UTC on 29 Sep 2015 (about 2 years ago)
First posted at 15:43 UTC on 29 Sep 2015 and last updated at 17:40 UTC on 29 Sep 2015.
Resolved – Global Issue: Imaging slow on BigV
Resolved at 16:35 UTC on 27 Sep 2015 (about 2 years ago)
First posted at 12:06 UTC on 27 Sep 2015 and last updated at 16:35 UTC on 27 Sep 2015.
Resolved – Global Issue: Advisory: Introduction of a new IP transit provider
Resolved at 19:16 UTC on 23 Sep 2015 (about 2 years ago)
First posted at 16:55 UTC on 23 Sep 2015 and last updated at 19:16 UTC on 23 Sep 2015.
Resolved – Meta Issue: Introduction of new Core Router
Resolved at 22:33 UTC on 21 Sep 2015 (about 2 years ago)
First posted at 17:17 UTC on 21 Sep 2015 and last updated at 22:33 UTC on 21 Sep 2015.
Resolved – Meta Issue: BigV head22 unexpected reboot
Resolved at 06:03 UTC on 19 Sep 2015 (about 2 years ago)
First posted at 04:25 UTC on 19 Sep 2015 and last updated at 06:03 UTC on 19 Sep 2015.
Resolved – Meta Issue: Maintenance to status.bytemark.org
Resolved at 11:11 UTC on 17 Sep 2015 (about 2 years ago)
First posted at 09:15 UTC on 17 Sep 2015 and last updated at 11:11 UTC on 17 Sep 2015.
Resolved – Meta Issue: BigV Manchester switch upgrade
Resolved at 13:29 UTC on 10 Sep 2015 (about 2 years ago)
First posted at 11:40 UTC on 10 Sep 2015 and last updated at 13:29 UTC on 10 Sep 2015.
Resolved – Meta Issue: Our mirror service will be undergoing maintainance on Tuesday 7th September.
Resolved at 12:46 UTC on 08 Sep 2015 (about 2 years ago)
First posted at 14:12 UTC on 05 Sep 2015 and last updated at 12:46 UTC on 08 Sep 2015.
Resolved – Meta Issue: BigV signup not working
Resolved at 18:20 UTC on 07 Sep 2015 (about 2 years ago)
First posted at 17:56 UTC on 07 Sep 2015 and last updated at 18:20 UTC on 07 Sep 2015.
Resolved – Global Issue: Reports of packet loss across Level3's network
Resolved at 23:35 UTC on 04 Sep 2015 (about 2 years ago)
First posted at 21:47 UTC on 04 Sep 2015 and last updated at 23:35 UTC on 04 Sep 2015.
Resolved – Global Issue: SSO password corruption
Resolved at 16:22 UTC on 03 Sep 2015 (about 2 years ago)
First posted at 16:01 UTC on 03 Sep 2015 and last updated at 16:22 UTC on 03 Sep 2015.
Resolved – Meta Issue: Internal database blip affecting some sites
Resolved at 22:15 UTC on 01 Sep 2015 (about 2 years ago)
First posted at 22:09 UTC on 01 Sep 2015 and last updated at 22:15 UTC on 01 Sep 2015.
Resolved – Global Issue: Can't create new VMs in York
Resolved at 16:09 UTC on 28 Aug 2015 (about 2 years ago)
First posted at 14:36 UTC on 28 Aug 2015 and last updated at 16:09 UTC on 28 Aug 2015.
Resolved – Meta Issue: Ticket System Scheduled Maintenance
Resolved at 20:32 UTC on 20 Aug 2015 (about 2 years ago)
First posted at 18:12 UTC on 20 Aug 2015 and last updated at 20:32 UTC on 20 Aug 2015.
Resolved – Meta Issue: Unavailable: New domain registrations & inbound transfers via panel
Resolved at 14:43 UTC on 13 Aug 2015 (over 2 years ago)
First posted at 11:24 UTC on 12 Aug 2015 and last updated at 14:43 UTC on 13 Aug 2015.
Resolved – Global Issue: Backup server Waltz out of service
Resolved at 10:59 UTC on 13 Aug 2015 (over 2 years ago)
First posted at 13:29 UTC on 12 Aug 2015 and last updated at 10:59 UTC on 13 Aug 2015.
Resolved – Dedicated - mirror: Mirror Server Down
Resolved at 23:54 UTC on 06 Aug 2015 (over 2 years ago)
First posted at 22:45 UTC on 06 Aug 2015 and last updated at 23:54 UTC on 06 Aug 2015.
Resolved – BigV - bigv1-tail4: BigV Archive Tail Unavailable
Resolved at 11:10 UTC on 01 Aug 2015 (over 2 years ago)
First posted at 10:57 UTC on 01 Aug 2015 and last updated at 11:10 UTC on 01 Aug 2015.
Resolved – Global Issue: Backup space waltz unavailable
Resolved at 22:46 UTC on 24 Jul 2015 (over 2 years ago)
First posted at 08:46 UTC on 24 Jul 2015 and last updated at 22:46 UTC on 24 Jul 2015.
Resolved – Global Issue: Some peering maintenance
Resolved at 02:55 UTC on 24 Jul 2015 (over 2 years ago)
First posted at 20:56 UTC on 23 Jul 2015 and last updated at 02:55 UTC on 24 Jul 2015.
Resolved – Global Issue: bigv1-head2 blip
Resolved at 22:11 UTC on 23 Jul 2015 (over 2 years ago)
First posted at 21:45 UTC on 23 Jul 2015 and last updated at 22:11 UTC on 23 Jul 2015.
Resolved – BigV - bigv1-head24: VMs Down/Restarted
Resolved at 11:46 UTC on 23 Jul 2015 (over 2 years ago)
First posted at 10:48 UTC on 23 Jul 2015 and last updated at 11:46 UTC on 23 Jul 2015.
Resolved – Global Issue: London peering offline
Resolved at 08:59 UTC on 05 Jul 2015 (over 2 years ago)
First posted at 01:05 UTC on 05 Jul 2015 and last updated at 08:59 UTC on 05 Jul 2015.
Resolved – Global Issue: A denial of service attack is affecting some connectivity.
Resolved at 23:57 UTC on 03 Jul 2015 (over 2 years ago)
First posted at 23:14 UTC on 03 Jul 2015 and last updated at 23:57 UTC on 03 Jul 2015.
Resolved – BigV - bigv1-head18: BigV head18 reboot (~ 1735 28 June 2015 UTC+1)
Resolved at 14:22 UTC on 29 Jun 2015 (over 2 years ago)
First posted at 14:22 UTC on 29 Jun 2015.
Resolved – Meta Issue: Fasthosts DNS (livedns.co.uk) is broken
Resolved at 19:02 UTC on 18 Jun 2015 (over 2 years ago)
First posted at 18:04 UTC on 18 Jun 2015 and last updated at 19:02 UTC on 18 Jun 2015.
Resolved – Global Issue: Packet Loss
Resolved at 18:02 UTC on 18 Jun 2015 (over 2 years ago)
First posted at 13:51 UTC on 18 Jun 2015 and last updated at 18:02 UTC on 18 Jun 2015.
Resolved – BigV - bigv1-head25: Instability on BigV head 25
Resolved at 22:06 UTC on 14 Jun 2015 (over 2 years ago)
First posted at 21:47 UTC on 14 Jun 2015 and last updated at 22:06 UTC on 14 Jun 2015.
Resolved – Meta Issue: Careers site unavailable
Resolved at 17:06 UTC on 03 Jun 2015 (over 2 years ago)
First posted at 17:04 UTC on 03 Jun 2015 and last updated at 17:06 UTC on 03 Jun 2015.
Resolved – Meta Issue: Bytemark internal database server fault: forum, support, fault alerting down from 0005-0915 on Sat 30th May
Resolved at 12:49 UTC on 03 Jun 2015 (over 2 years ago)
First posted at 12:49 UTC on 03 Jun 2015.
Resolved – Meta Issue: IX Manchester port upgrade
Resolved at 14:47 UTC on 02 Jun 2015 (over 2 years ago)
First posted at 09:47 UTC on 02 Jun 2015 and last updated at 14:47 UTC on 02 Jun 2015.
Resolved – Global Issue: Level3 / Verizon routing issue causing reachability problems for Bytemark network
Resolved at 01:15 UTC on 29 May 2015 (over 2 years ago)
First posted at 00:31 UTC on 29 May 2015 and last updated at 01:15 UTC on 29 May 2015.
Resolved – BigV - bigv1-brain2: BigV UK0 brain down
Resolved at 16:41 UTC on 28 May 2015 (over 2 years ago)
First posted at 16:27 UTC on 28 May 2015 and last updated at 16:41 UTC on 28 May 2015.
Resolved – Router - fw2.man: DoS attack
Resolved at 16:40 UTC on 28 May 2015 (over 2 years ago)
First posted at 16:40 UTC on 28 May 2015.
Resolved – Global Issue: Some trouble at LINX
Resolved at 14:25 UTC on 27 May 2015 (over 2 years ago)
First posted at 10:00 UTC on 27 May 2015 and last updated at 14:25 UTC on 27 May 2015.
Resolved – Meta Issue: Database server restart
Resolved at 17:40 UTC on 25 May 2015 (over 2 years ago)
First posted at 17:20 UTC on 25 May 2015 and last updated at 17:40 UTC on 25 May 2015.
Resolved – Global Issue: Emergency maintenance on legacy virtual machines 15 May 2015 0700 UTC+1
Resolved at 16:08 UTC on 15 May 2015 (over 2 years ago)
First posted at 16:08 UTC on 15 May 2015.
Resolved – Meta Issue: CVE-2015-3456 aka "Venom": new vulnerability affecting virtual machine platforms (BigV and legacy VMs)
Resolved at 11:18 UTC on 15 May 2015 (over 2 years ago)
First posted at 14:31 UTC on 13 May 2015 and last updated at 11:18 UTC on 15 May 2015.
Resolved – VirtualMachineHost - host08.vmh: Networking Issues
Resolved at 15:38 UTC on 14 May 2015 (over 2 years ago)
First posted at 15:37 UTC on 14 May 2015 and last updated at 15:38 UTC on 14 May 2015.
Resolved – Global Issue: Sustained DDoS Attack
Resolved at 09:28 UTC on 08 May 2015 (over 2 years ago)
First posted at 02:18 UTC on 08 May 2015 and last updated at 09:28 UTC on 08 May 2015.
Resolved – Meta Issue: Bytemark sites down
Resolved at 15:28 UTC on 30 Apr 2015 (over 2 years ago)
First posted at 15:26 UTC on 30 Apr 2015 and last updated at 15:28 UTC on 30 Apr 2015.
Resolved – Global Issue: Level3 IP transit interruption
Resolved at 12:27 UTC on 24 Apr 2015 (over 2 years ago)
First posted at 13:09 UTC on 23 Apr 2015 and last updated at 12:27 UTC on 24 Apr 2015.
Resolved – Global Issue: DNS Resolutions Failing
Resolved at 10:53 UTC on 24 Apr 2015 (over 2 years ago)
First posted at 10:53 UTC on 24 Apr 2015 and last updated at 10:53 UTC on 24 Apr 2015.
Resolved – Global Issue: panel-beta not allowing logins
Resolved at 13:51 UTC on 16 Apr 2015 (over 2 years ago)
First posted at 13:41 UTC on 16 Apr 2015 and last updated at 13:51 UTC on 16 Apr 2015.